A mixture of CH4 and H2O is passed over a nickel catalyst at 1000K. The emerging gas is collected in a 5.00L flask and is found to contain 8.62g of CO, 2.60g of H2, 43.0g of CH4 and 48.4g of H2O. Assuming that equilibrium has been reached, calculate Kc and Kp for the reaction.


1
Expert's answer
2021-06-07T03:16:07-0400

According to that Kc is an equilibrium constant in terms of molar concentrations.

and Kc = [C]^c *[D]^d / [A]^a * [B]^b >>>> (1)

in the general reaction:

aA + bB ↔ cC + dD 

and, from our balanced equation:

CH4 + H2O ⇔ Co + 3H2 >>> (2)

So, we need to calculate the concentrations (molarity) of the products and reactants:

the Molarity of CH4 = no. of moles/volume (L)

 and no. of moles = weigh / Molecular weight = 42.3 / 16 = 2.643 moles

so the molarity of CH4 = 2.643 / 5 = 0.528 molar

the molarity of H20 = (49.2 / 18) / 5 = 0.546 molar

the molarity of CO = (8.32/28) / 5 = 0.059 molar

the molarity of H2 = (2.63 / 2) / 5 = 0.263 molar 

By substitution in (1) according to (2);

∴ Kc = [0.059]*[0.263]^3 / ( [0.528]*[0.546]) = 3.7 * 10 ^-3 >>>> (3)

Kp = Kc (RT)^(Δn) >>> (4)

where R is the gas constant = 0.0821,

and Δn is the change in moles in gas= (3(H2) + 1 (CO) - (1 H2O + 1 CH4) = 2

by substition in (4):

∴ Kp = 3.7*10^-3 (0.0821* 1000)^2= 24.939